OpenStudy (zuha):

Simplify the expression. (9+4i)^2

OpenStudy (mathmale):

Try expanding (9+4x)^2 first. Then tackle (9+4i)^2. Note: "i" is the imaginary operator: i=sqrt(-1) i^2= -1 i^3= -i i^4=1

OpenStudy (zuha):

@mathmale 81+16x?

pooja195 (pooja195):

"^2" basically means it's written twice \[\huge~\rm (9+4i)^2--> (9+4i)(9+4i)\]

pooja195 (pooja195):

Now just use the Foil method Can you tell me what we get once we do that?

OpenStudy (zuha):

@pooja195 81+36i+36i+4i^2

OpenStudy (zuha):

So basically 81+72i+4i^2

pooja195 (pooja195):

"^2" basically means it's written twice \[\huge~\rm (9+4i)^2--> (9+4i)(9+4i)\] 81+36i+36i+16i^2

OpenStudy (zuha):

Ah makes sense, so 81+72i+16i^2?

pooja195 (pooja195):

Keep in mind that i^2=-1

pooja195 (pooja195):

"^2" basically means it's written twice \[\huge~\rm (9+4i)^2--> (9+4i)(9+4i)\] 81+36i+36i+16(-1) Combine the like terms :) we got: 81+36i+36i-16

pooja195 (pooja195):

Are you stuck?

OpenStudy (zuha):

Ah so the i would be removed from 16 as there is as a square, correct?

OpenStudy (zuha):

So would it be 81+72i+16?

pooja195 (pooja195):

not quite 72i is correct but remember we said i^2=-1 which makes 16i^2=-16 Which would mean we subtract 81-16=?

OpenStudy (zuha):

65!

pooja195 (pooja195):

Yes! :D so whats the answer? :O

OpenStudy (zuha):

65+72i

OpenStudy (zuha):

That's perfect, thank you!!!

pooja195 (pooja195):

Perfect <3 You're welcome ^_^